OpenStudy (anonymous):

What is the approximate distance from (-2,3) to (4,1)? 5 units 6 units 8 units 9 units

OpenStudy (anonymous):

6

OpenStudy (anonymous):

6 units is the answer

OpenStudy (anonymous):

\[distance=\sqrt{(x2-x1)^2+(y2-y1)^2}\] \[\sqrt{(1-3)^2+(4-(-2)}\] \[\sqrt{(-2)^2+(6)^2}\] \[\sqrt{4+36}\] \[\sqrt{40}\] \[6.3=6units\]

OpenStudy (amistre64):

one way to determine the distance is to create a rt triangle and use the pythag thrm on it

OpenStudy (amistre64):

if we move from the lowest point up to the origin, we can get out lengths for x and y\[\] (-2+2,3-3) to (4+2,1-3)\[\] (0,0) to (6,-2)\[\] this tells us x and y are 6 and -2, regardless of which is which; we can use them in the pythag

OpenStudy (amistre64):



OpenStudy (anonymous):

Thanks guyss :D i think this is my last question 4 a while :)

OpenStudy (amistre64):

6^2 + 2^2 = d^2\[\] 36 + 4 = d^2\[\] 40 = d^2\[\] which option when squared is closest to 40?

OpenStudy (anonymous):

6 :D

OpenStudy (amistre64):

id say 6 as well :) just to conform with the masses lol